# Deep-Learning-Driven Prefetching for Far Memory

### v1

@@ 4.1. End-to-End Application Performance
Application performance. Figures 11, 11, 11, and 11 present the end-to-end application performance of MCF, XGBoost, PageRank, and SSSP, respectively. For each set of experiments, we change the application server’s local memory size from 30% to 90% of the total application memory size (X axis) and measure the total application execution time (Y axis). For each result, we normalize the application execution time against that of running at full local-memory capacity, and higher Y-axis values are better.
@@
Across all four workloads and input settings, FarSight consistently outperforms FastSwap and Hermit by up to 3.6 times and 2.6 times, respectively. Comparing across local-memory size settings, FarSight achieves greater improvements when the local memory size is small—an environment especially challenging but useful for far-memory systems. A smaller local memory size increases the likelihood of missed accesses, amplifying the performance impact of an effective prefetching policy.
@@
Figure 16. Handling Input Variance in MCF. Three inputs used for MCF (220 MB, 240 MB, and 390 MB) tested on the same trained model.
Figure 17. Input Variation in PageRank. Three graphs used for PageRank (2.2 GB, 3.9 GB, and 8.2 GB) tested on the same trained model.
@@
Figure 18. Performance Breakdown of FarSight using MCF with $30\%$ local memory. Each bar adds one of FarSight’s techniques at a time.
FarSight outperforms the rule-based far-memory baseline systems because it is able to make more future access predictions and thus issue more prefetches. Moreover, the prefetches have high hit rates (i.e., highly accurate). We will present further analysis of these effects with additional evaluations shortly.
Among the four workloads, FarSight performs best on XGBoost, where our DL-based prediction framework effectively captures the tree traversal pattern; however, the heuristic-based strategies used in the baseline systems fail to do so. MCF is the hardest to prefetch among the four workloads because of its irregular graph-processing patterns, as can be seen from the overall lower performance when local memory is small. Nonetheless, FarSight manages to improve its performance by up to 2.6 and 3.6 times over Hermit and FastSwap. FarSight’s improvement is relatively lower with the two GAP workloads. Both GAP PageRank and SSSP represent the input graphs in matrices form and perform traversals on them: PageRank iteratively accesses the graph across multiple rounds, while SSSP performs a one-time traversal. Because of the compact matrix representation of graph, these workloads exhibit more sequential patterns that can be captured by the baselines.
@@
Figure 19. Performance Breakdown of FarSight using MCF with $50\%$ local memory. Each bar adds one of FarSight’s techniques at a time.
Figure 20. Effect of Vocabulary Size on End to End Performance.
Figure 21. Effect of Vocabulary Size on Prefetch Hit Rate.
Prefetch hits and rates. To understand the effectiveness of FarSight’s DL-based prefetching, we analyze the total number of prefetches that are accessed by applications (i.e., prefetch hit count) and the ratio of them to the total amount of prefetches (i.e., prefetch hit rate). Figure 15 and Figure 15 present the number of prefetch hits with different local memory ratios using the MCF and XGBoost workloads, respectively. Figure 15 and Figure 15 present the corresponding prefetch rates. We do not include Hermit in these figures, as their prefetch events are stopped when being determined as unuseful, thus having even fewer prefetch hits than FastSwap.
For MCF, FarSight issues 15.3×15.3$\times$15.3 × more prefetch requests and achieves 13.2×13.2$\times$13.2 × more prefetch hits on average compared to FastSwap. This shows that FarSight can uncover more application memory-access patterns and turn them into fruitful prefetches. Similarly, FarSight issues significantly more prefetch requests than FastSwap for XGBoost.
Interestingly, FarSight’s prefetch hit rate is lower than FastSwap’s, although still high, ranging from 67% to 97% for the two workloads. FastSwap’s high prefetch hit rate is because of its conservative prefetching policy—only when a sequential access pattern is detected will it perform prefetching. As sequential patterns are regular, FastSwap’s hit rate is consistently high. However, this conservative behavior results in FastSwap’s overall lower prefetch hit count and worse end-to-end application performance.
4.2. Deep Dive
We now perform a deep dive to understand and evaluate FarSight’s performance benefits.
Generalization to different inputs. To demonstrate our model’s ability to capture program semantics across program inputs, we evaluate the end-to-end application performance using multiple inputs with different sizes with the same model trained by another input. Figure 18 and Figure 18 show the results of these experiments. To challenge FarSight  we train the model on a small dataset and with larger, unseen inputs.
For MCF, we trained the DL model using a graph with 5K nodes and 50K edges. We test it with three different graphs, each with a distinct structure that is different from the training graph and contains between 20K and 40K nodes and over 200K edges. For PageRank, the model is trained using a graph with 10k nodes and 100k edges, and is tested on graphs with 4M, 8M, and 16M nodes, respectively. Both applications were tested under $30\%$ and $50\%$ local memory ratios.
FarSight demonstrates strong generalization capability with one-time offline training, effectively adapting to a wide range of input structures and runtime environments without the need for retraining or fine-tuning. In the case of MCF, this result in an average performance improvement of 2.3×2.3$\times$2.3 × over FastSwap, while in PageRank, we observe a 1.9×1.9$\times$1.9 × gain. FarSight is able to generalize across inputs and adapt to larger inputs with models trained on smaller ones, thanks to our memory-access behavior and address layout decoupling. These results demonstrate that effective models can be trained using short sequences of memory accesses as training data points, provided they reflect the application’s core memory-access patterns. This approach can substantially reduce both the training data collection overhead and the computational cost of model training.
Ablation study. To understand the performance benefits of each individual technique used in FarSight, we evaluate the system incrementally by adding one technique at a time, as shown in Figure 18 and Figure 21 using the MCF workload with 30% and 50% local memory. We present results under two different memory ratios to highlight how each technique performs under varying levels of memory pressure.
The leftmost bar in each figure represents a baseline configuration—the vanilla Linux setup without any prefetching. We then add FarSight’s DL-based prediction and trigger the prediction and prefetching synchronously on every page fault, including both faults that result in a local-memory miss and faults that hit the swap cache. Performing this basic DL-based prefecthing has a small improvement over the baseline.
@@ 4.2. Deep Dive
We then change the prediction and prefetching to only be triggered on page misses, as shown by the third bars. Although this gives FarSight less chance to perform prediction, we can significantly avoid the runtime performance overhead that would otherwise be incurred for prediction on page hits. This is because a page fault that hits the swap cache will issue no far-memory I/O, leaving insufficient time for FarSight to finish its prediction.
So far, the prefetching FarSight performs is synchronous—application threads wait for it to finish. We then change the system to perform prefetching asynchronously, as shown in the fourth bars. FarSight optimizes latency by allowing the system to return to userspace as soon as the on-demand page is ready, rather than waiting for all prefetched pages to complete.
Next, we add the technique of swappable and indirected future maps. By proactively evicting cold future maps, FarSight frees up memory for user applications, leading to further performance improvements.
The final technique incorporated into the full FarSight design focuses on improving the timeliness of prefetching to avoid late prefetch—–a scenario where a page fault still incurs I/O latency even though the page was prefetched. By predicting and issuing prefetches earlier, FarSight ensures that prefetched pages are more likely to arrive before they are actually needed. As the latency results indicate, this strategy effectively reduces blocking and provides the final boost in system performance.
@@
Sensitivity Tests By default, FarSight uses a vocabulary and future-map size of 64 to strike a balance between memory overhead and prediction accuracy. To evaluate how FarSight performs under different vocabulary size, we change $K$ from 8 to 64. Figure 21 and Figure 21. Overall, FarSight’s benefits prevail across $K$ values compared to FastSwap and Hermit, showing its robustness to sensitivity. FarSight’s performance degrades slightly when $K$ is smaller than 32. This is because a small future map size cannot capture the possible outcomes of application memory accesses. For example, consider an application with a recurring access pattern involving 1q frequently accessed pointers in an alternating sequence. If the future map size is limited to just 8 entries, the prediction pointers will overwrite each other, leading to reduced prefetch accuracy and a lower hit rate. This is evidenced by the decreased prefetch hit rates in Figure 21. Meanwhile, future map sizes larger than 64 add runtime performance overhead both becaseue future maps need to be swapped more frequently and because they can be less cached in CPU cache. Thus, we set the vocabulary size to be 64 by default, which works well for many typical applications.
5. Related Work
This section discusses related works in ML-based CPU cache and far-memory prefetching.
5.1. ML-Based CPU Cache Prefetchers
Prior research works have explored using ML techniques for local server prefetch predictions at the micro-architecture level by prefetching data from memory into CPU cache. However, because of their performance and/or accuracy issues, they have only been realized in simulation or for offline trace analysis. For example, Peled et al. have proposed reinforcement-learning-based (peled2015semantic, ) and regression-based (peled2019neural, ) approaches for memory prefetching. The former sets up the prefetching prediction as a classification problem and can only accommodate four possible address offsets for each prediction. The latter has accuracy issues as a regression model aims to be close to the ground truth, but correct prefetch requires the exact truth.
Another series of research works use LSTM to predict memory access sequences in the form of memory addresses (voyager, ; pmlr-v80-hashemi18a, ; srivastava2019predicting, ). The major problem with these solutions is their large vocabulary size, which results in slower prediction and low model accuracy (especially small ones). For example, no sequence-to-sequence models can handle a vocabulary size of $2^{48}$ — the possible addresses in a 64-bit address space for the current Intel CPU. To reduce the vocabulary size, Hashemi et al. (pmlr-v80-hashemi18a, ) limit the prediction to address deltas within a spatial region. However, this approach prevents predictions of data accesses that are distant from each other. Voyager (voyager, ) decouples memory pages and offsets within a page and only covers addresses that an application uses. A major issue with this approach is that a trained model with one application execution cannot be used by another execution, as different inputs or address randomization can all result in different sets of used addresses. As a result, Voyager requires an online model training every million memory accesses and “does not yet make neural models practical” (voyager, ).
More recently, Neural Network and Transformer-based models have also been applied to CPU cache prefetching (duong2024twilight, ; zhang2023dart, ). Twilight and T-LITE (duong2024twilight, ) use the combination of a customized two-layer neural-network model, clustering, and frequency-based history table for CPU cache prefetching. DART (zhang2023dart, ) distills a transformer model and then transforms the distilled model into a hierarchy of table lookups to reduce runtime performance overhead. Although these works have shown their CPU cache prefetch effectiveness through simulation, their training and prediction processes are complex and lack generalization or consistent accuracy.
Overall, unlike FarSight, these micro-architecture level CPU cache prefetching systems are unfit for real deployment in far-memory runtime systems, as they cannot achieve high prefetch accuracy, low overhead, and application generalization simultaneously. FarSight demonstrates the unique challenges and opportunities of applying ML techniques to software system problems, examplified by the challenging far-memory prefetching problem.
5.2. Far-Memory Prefetching
Existing far memory systems have explored various approaches for improving swap and data prefetching performance. On the mechanism side, Fastswap (fastswap, ) optimizes the swap datapath, and Hermit (hermit, ) further proposes asynchronous, proactive swap-out techniques. These systems improve execution efficiency, but still rely on the default Linux prefetching policy, which only follows simple rules and often fails to trigger under many real-world applications. Leap (almaruf2020leap, ) enhances the Linux prefetcher by attempting to match application page access patterns to a set of predefined trends, such as sequential and strided patterns. However, it struggles with complex, runtime-dependent application memory access patterns.
Different from swap-based far-memory systems, Mira (mira, ) is a fine-grained far-memory system that incorporates static program analysis, compiler optimization, and runtime profiling. Mira captures static program semantics and inserts prefetching operations with its compiler transformation. However, its performance degrades for applications that depend heavily on runtime information, such as graph-based workloads. Moreover, for applications to use Mira and other library-level far-memory solutions (aifm-osdi20, ), applications need to be ported with non-trivial manual effort.
Unlike FarSight, these prior systems fall short in far-memory prefetching scenarios, where access patterns are constructed dynamically at runtime, making them difficult to capture with pre-defined heuristics and highly sensitive to variations across application inputs.
6. Conclusion
@@
We presented FarSight, a DL-based far-memory prefetching system in the Linux kernel. FarSight’s core idea is to decouple the learning of application semantics from the runtime capturing of memory accesses. By doing so and with our set of optimization techniques, FarSight achieves overall application performance benefits over two recent far-memory systems, by up to 3.6 times and 2.6 times. FarSight demonstrates the feasibility of deploying modern ML techniques to solve performance-critical problems in complex runtime systems. Future systems researchers and practitioners could leverage lessons we learned and building blocks of FarSight’s DL model, prediction problem presentation, and system-integration mechanisms.
Acknowledgements.
@@
We would like to thank Ryan Lee, Geoff Voelker, Zijian He, Vikranth Srivatsa, and Reyna Abhyankar for their valuable contributions and feedback on this paper. This material is based upon work supported by funding from PRISM center (part of SRC’s JUMP 2.0) and gifts from AWS, Google, and Meta. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of these institutions.

### v2

@@ 4.1 End-to-End Application Performance
Application performance. Figures 7, 7, 7, and 7 present the end-to-end application performance of MCF, XGBoost, PageRank, and SSSP, respectively. For each set of experiments, we vary the application server’s local memory size between 30% and 90% of the total application memory size (X axis) and measure the total application execution time (Y axis). For each result, we normalize the application execution time against that of running at full local-memory capacity, and higher Y-axis values are better.
Figure 8: MCF F1 score.
Figure 9: MCF recall.
Figure 10: MCF precision.
Figure 11: XGBoost F1 score
Figure 12: XGBoost recall
Figure 13: XGBoost precision
Figure 14: Shortest path F1 score
Figure 15: Shortest path recall
Figure 16: Shortest path precision
Figure 17: Page rank F1 score
Figure 18: Page rank recall
Figure 19: Page rank precision
@@
FarSight consistently delivers superior performance across all four workloads, achieving up to 3.6 times and 2.6 times speedups over FastSwap and Hermit, and outperforming Twilight on MCF by up to 3.9 times. Comparing across local-memory size settings, FarSight achieves greater improvements when the local memory size is small—an environment especially challenging but useful for far-memory systems. A smaller local memory size increases the likelihood of missed accesses, amplifying the performance impact of an effective prefetching policy. Among the four workloads, FarSight performs best on MCF, where our DL-based prediction framework effectively captures the graph traversal pattern.
Prefetch precision and recall. To understand the effectiveness of FarSight’s DL-based prefetching, we analyze the prefetch effectiveness by measuring precision, recall, and F1 score of its prediction and the baselines’. In the far-memory scenario, we define precision as the fraction of correctly predicted pages that are accessed by the application in local memory after prefetching (i.e., true positive) over the total number of predictions (i.e., prefetches). We define recall as the fraction of correctly predicted pages over the number of misses (page faults) plus the number of correctly predicted pages (i.e., false negative ++ true positive). Figures 1019 plot the F1 score, recall, and precision of MCF, XGBoost, Page Rank, and Shortest Path workload.
Overall, FarSight’s F1 score and recall are higher than all the baselines across settings and workloads, explaining its superior end-to-end application performance. FarSight eliminates 11% to 17% on-demand page faults compared to Hermit, 37% to 63% page faults compared to FastSwap and over 50% page faults compared to Twilight.
FarSight’s precision is much higher than Twilight and on par with FastSwap and Hermit when local memory is large. When local memory is small, its precision is lower than FastSwap and Hermit. This is because FastSwap and Hermit rely on Linux’s sequential prefetcher; only when a sequential pattern is detected do these systems perform prefetching, resulting in their high accuracy but low recall. FarSight takes a more agressive approach by trying to predict more complex access patterns, explaining its higher recall and relatively lower precision.
Comparison to Twilight. Twilight has the lowest precision and recall and performs the worst among all systems for several reasons. First, it involves two sub-predictions, one to determine the behavioral cluster an address belongs to and one to predict the subsequent for an address within a cluster. This process compounds inaccuracy in each sub-prediction, resulting in Twilight’s overall lower accuracy. Second, it predicts the immediate next access. In far-memory systems, fetching a page requires one network round-trip. By the time the predicted page is fetched, the application usually already has the need to access it, still causing a page fault and explaining Twilight’s low recall.
4.2 Performance Deep Dive
Generalization to different inputs. To demonstrate FarSight’s ability to generalize across inputs, we evaluate end-to-end application performance using multiple inputs of varying sizes. Figure 23 presents the results of MCF under $30\%$ and $50\%$ local memory ratios. The model is trained using a graph with 5K nodes and 50K edges and tested using three different graphs containing 20K to 40K nodes and over 200K edges. FarSight demonstrates strong generalization capability with one-time offline training, with an average performance improvement of $2.3\times$ over FastSwap across inputs. FarSight is able to generalize across inputs and adapt to larger inputs with models trained on smaller ones, thanks to our memory-access behavior and address layout decoupling.
@@ 4.2 Performance Deep Dive
Figure 20: Handling input variance in MCF. Three inputs used for MCF (220 MB, 240 MB, and 390 MB) tested on the same trained model.
@@
Figure 21: Performance breakdown of FarSight using MCF with $30\%$ local memory. Each bar adds one of FarSight’s techniques at a time.
@@
Figure 22: Performance breakdown of FarSight using MCF with $50\%$ local memory. Each bar adds one of FarSight’s techniques at a time.
Figure 23: Effect of vocabulary size on end to end performance.
@@
Ablation study. To understand the performance benefits of each individual technique used in FarSight, we evaluate the system incrementally by adding one technique at a time. Figure 2323 shows the breakdown with the 30% and 50% local memory configuration. The leftmost bar represents a baseline configuration—the vanilla Linux setup without any prefetching. First, we add FarSight’s model prediction but perform the prediction and prefetching synchronously on every page fault, which improves the baseline but not significantly. We then change the prediction and prefetching to only be triggered on page misses, as shown by the third bar. Although this gives FarSight less chance to perform prediction, we can significantly avoid the runtime performance overhead that would otherwise be incurred for prediction on page hits, resulting in improved performance.
So far, the prefetching FarSight performs is synchronous—application threads wait for it to finish. The fourth bar demonstrates the performance gain when performing prefetching asynchronously and allowing more flexible memory swapping, which improves the application performance further. The final technique incorporated into the full FarSight design incorporating the optimizations of look-ahead, batched prefetching. By predicting and issuing prefetches earlier, FarSight ensures that prefetched pages are more likely to arrive before they are actually needed.
@@
Sensitivity Tests By default, FarSight uses a vocabulary and future-map size of 64 to strike a balance between memory overhead and prediction accuracy. To evaluate how FarSight performs under different vocabulary size, we vary $K$ between 8 and 64. As shown in Figure 23, FarSight demonstrates its robustness to various vocabulary size and consistently outperforms FastSwap and Hermit. FarSight’s performance degrades slightly when $K$ is smaller than 32. This is because a small future map size cannot capture the possible outcomes of application memory accesses. For example, in an application with a recurring access pattern involving 16 frequently accessed pointers in an alternating sequence, a future map size of 8 entries will cause prediction pointers to overwrite each other, reducing prefetch accuracy and lowering the hit rate. Meanwhile, future map sizes larger than 64 add runtime performance overhead as they are less likely to be cached in CPU. Thus, we set the vocabulary size to be 64 by default, which works well for many typical applications.
5 Conclusion
@@
We presented FarSight, a DL-based far-memory prefetching system with a core idea of decoupling the learning of application semantics from the runtime capturing of memory accesses. By doing so and with our set of optimization techniques, FarSight achieves overall application performance benefits over two recent far-memory systems, by up to 3.6 times and 2.6 times, and a SOTA CPU prefetcher by 3.9 times. FarSight demonstrates the feasibility of deploying modern ML techniques to solve performance-critical problems in complex runtime systems. Future researchers and practitioners could leverage lessons we learned and building blocks of FarSight’s DL model, prediction problem presentation, and system-integration mechanisms.
6 Acknowledgment
@@
We would like to thank Ryan Lee, Geoff Voelker, Zijian He, Vikranth Srivatsa, and Reyna Abhyankar for their valuable contributions and feedback on this paper. This material is based upon work supported by funding from PRISM center (part of SRC’s JUMP 2.0), NSF award 2403253, and gifts from AWS, Google, and Meta. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of these institutions.